A square lobby is to be tiled at a cost 160 per square meter. If the lobby were 2 m narrower and 2 m longer the cost of the tiling would be 35,360. Find the dimensions of the lobby.
Answer by edjones(8007)

s=one side of the square lobby.
35360/160=221 m^2
(s+2)(s-2)=221
s^2-4=221
s^2=225
s=15 m
.
Ed